# HWA-UNETR: Hierarchical Window Aggregate UNETR for 3D Multimodal Gastric Lesion Segmentation

### v1

@@ 3.4 Tri-orientated Fusion Mamba Block
where $X_{mo}$ denotes the output of TFM Block; $S$ denotes the softmax function, $Ma$ refers to the Mamba module that models global information within the sequence, $X_{so.f}$, $X_{so.r}$, and $X_{so.s}$ denote the forward, reverse, and inter-slice direction feature maps, respectively.
@@ 4.3 Performance
Visual Comparisons : Fig. 3 provides a more intuitive comparison of the segmentation results obtained by different methods. On the GCM 2025 dataset, our HWA-UNETR accurately detects the boundaries of tumor regions in each modality image, demonstrating better consistency compared to other SOTA methods.
@@ 5.0.1 Acknowledgements
This study was funded by the Key-Area Research and Development Program of Guangzhou City (2023B01J1001), the National Key Research and Development Program of China (2022ZD0117700), the National Natural Science Foundation of China (U21A20520, 62325204), the Key-Area Research and Development Program of Guangzhou City (2023B01J1001), and the Hainan University High-level Talent Research Launch Fund (XJ2400012551).

### v2

@@ 3.4 Tri-orientated Fusion Mamba Block
where $X_{mo}$ denotes the output of TFM Block; $S$ denotes the softmax function [7], $Ma$ refers to the Mamba module that models global information within the sequence, $X_{so.f}$, $X_{so.r}$, and $X_{so.s}$ denote the forward, reverse, and inter-slice direction feature maps, respectively.
@@ 4.3 Performance
Visual Comparisons : Fig. 3 provides an intuitive comparison of the segmentation results obtained by different methods. On the GCM 2025 dataset, our HWA-UNETR accurately detects the boundaries of tumor regions in each modality image, demonstrating better consistency compared to other SOTA methods.
@@ 5.0.1 Acknowledgements
This work was supported in part by the National Key Research and Development Program of China (2022ZD0117700), the National Natural Science Foundation of China (U21A20520, 62325204), the Key-Area Research and Development Program of Guangzhou City (2023B01J1001), and the Hainan University High-level Talent Research Launch Fund (XJ2400012551).